OpenStudy (usukidoll):

Is it possible to find the equilibrium of these equations Y= X X = e^{-bY/X} b is a constant... setting both equations to 0, we have 0 = X 0 = e^{-bY/X}

OpenStudy (usukidoll):

now can we really find an equilibrium if we set these equations to 0 or do we have to substitute y = x into the second equation and then take the natural log.

OpenStudy (usukidoll):

@rational

OpenStudy (usukidoll):

the thing is .... if -bY/X is a regular fraction.. it's undefined unless y = 0 but since there is an exponential if y =0 and x = 0 we will only have the first equation satisfied, but the second equation will be 0=1 which is not true. Maybe taking the natural log of the second equation will work but ln 0 is also undefined so we can't do that. The same result happens when you take a natural log of a negative number...undefined.

OpenStudy (usukidoll):

these two equations are actually difference equations \[Y_{t+1}=X_t, x_{t+1}=e^{\frac{-bY_t}{x_t}}\]

OpenStudy (usukidoll):

subsitution values were given from this http://i.stack.imgur.com/MaKfP.jpg

OpenStudy (usukidoll):

I am on 2d. I know what 2c is already... I just need to find the equilibrium of those equations.

OpenStudy (usukidoll):

question should be does the equilibrium really exist with this? but there has to be one... otherwise what's the point in asking 2e if there's no equilibrium which is what x and y values make the equation 0 besides (0,0) which isn't working here

OpenStudy (usukidoll):

@ganeshie8